OpenStudy (anonymous):

the function H(x) is described below. H(x) = 3x - 4 What is the value of H(6)? A. 5 B. 6 C. 14

OpenStudy (anonymous):

To solve this, all you need to is replace every x in the equation given with the value of 6!

OpenStudy (anonymous):

H(6) means solve for the system H(x) for when x = 6.

OpenStudy (anonymous):

okay thanks.

OpenStudy (anonymous):

The answer is C, if you need to check your answer :) Good luck on your math endeavors! H(6) = 3(6) - 4 H(6) = 18 - 4 H(6) = 14

OpenStudy (anonymous):

thanks <3